OpenStudy (anonymous):

find the zeros of g(x)=4x^2-x+5 by using the quadratic formula

OpenStudy (aravindg):

do you know the quadratic formula ?

OpenStudy (abb0t):

\(\large \frac{ -b \pm \sqrt{b^2-4ac} }{ 2a }\), when you have: \(ax^2+bx+c\)

OpenStudy (abb0t):

\(ax^2+bx+c=0\)**